# Peak finding algorithm for cluster counting with domain adaptation

### 26 Feb 2024 v1

@@ 1 Introduction
One major challenge of the cluster counting is the efficient reconstruction algorithm to recover the clusters signals (in the form of peaks) in the waveform from a DC measurement. These cluster signals tend to be stacked together in noisy environments (Fig. 1). Traditional methods often fall short of achieving the necessary efficiency due to sub-optimal information utilization. For instance, in the derivative-based method (section A.1), the peaks are detected by analyzing derivatives and imposing threshold criteria. However, only the rising edges of the single pulses, instead of the full pulse shape, are effectively used. Machine learning (ML) algorithm, which are designed to harness large datasets to reduce complexity and find new features in data, are the state-of-the-art in PID.
@@
Traditional ML relies on fully labeled samples for supervised learning. Such method is well-suited for Monte Carlo (MC) simulated samples, where the MC samples are generated through first-principle simulation, ensuring perfect labels. However, when applying ML algorithm to real experimental data samples, it is usually difficult to apply fully supervised learning because of the following reasons. The first one is that discrepancies exist between the simulation and the experimental data, which leads to a degraded performance of the supervised model trained solely on simulated sample. In some cases, these models may even fail to predict experimental data accurately. The other reason is that the experimental data often lacks precise labels for fully supervised learning.
Domain adaptation (DA) aims at alleviating this issue by transferring knowledge between domains pan2009survey . In our context, the MC simulated sample serves as the source domain, while the experimental data represents target domain. A principled approach to solving domain adaptation involves aligning the source domain distribution with that of the target domain. This alignment allows us to leverage labeled data from the source domain to train a classifier in the aligned domain, which can then be applied to the target domain. In reference flamary2016optimal , Optimal Transport (OT) is exploited to align the source and target domains. One important property of OT is the geometry sensitivity, which facilitates distance computation between probability distributions. By relaxing the invariant requirement on conditional distribution, reference courty2017joint proposed the Joint Distribution Optimal Transport (JDOT) , in which both marginal feature and conditional distributions are adapted by minimizing a global divergence between them. A further improvement on the path is to integrate the deep neural network and enable large scale learning damodaran2018deepjdot .
In the context of these studies, models are trained with a target sample without any labels, which is also known as the unsupervised domain adaptation. In our research, we build upon the work presented in reference damodaran2018deepjdot and create a semi-supervised version of the domain adaptation model. The model will be employed for analyzing data collected during the drift-tube-prototype experiment conducted at CERN cern .
@@ 2 Data Samples for Cluster Counting
The major mechanism underlying for particle identification in gaseous detector is the ionization of matters by charged particles. For DC, each electron produced through ionization generates a distinct peak in the readout waveform. If the electronics is sufficient fast, these peaks become quantifiable. The reconstruction algorithm plays a crucial role in identifying and associating these peaks with individual electrons. The waveforms can be collected either from test beam experiment or from simulations. In the subsequent sections, we delve into the specifics of the collected data samples.
@@ 4.1 Experiment with Pseudo Data
Table 1: Area under curve (AUC) and partial-AUC results for models. The partial-AUC’s are the normalized AUC values with false positive rates (FPR) that less than 0.1.
@@ 5 Summary
Cluster counting in drift chamber represents a significant advacement in PID techniques for particle physics experiments. The reconstruction algorithm is a critical challenge in this context. Our semi-supervised DA algorithm achieves state-of-the-art performance for peak finding in real experimental data. Importantly, the algorithm’s principles can be extended to other DA problems that exhibit geometric sensitivity. Looking ahead, as more test beam data becomes available across a wider range of beam momenta, our algorithm will play a crucial role in data analysis and may be implemented in FPGA systems for feasibility studies related to cluster counting.
@@ Acknowledgements
This research was funded by National Natural Science Foundation of China (NSFC) under Contract Nos. 12275296, Institute of High Energy Physics (Chinese Academy of Sciences) Innovative Project on Sciences and Technologies under Contracts Nos. E3545BU210 and E25456U210.
@@ A.1 Derivative-Based Algorithm
The principle of a derivative-based algorithm is intuitive. If a peak has a fast rising edge, the derivative should be larger than some threshold, and one can detect peaks by requiring threshold passing. The pseudo code of the algorithm is summarized in Algorithm 1. To exclude noise in the baseline, a threshold $T_{1}$ of approximately three times the RMS of the average noise amplitudes is initially required. Higher order derivatives are efficient for hidden-peak detection, but more sensitive to noise. Therefore, a second-order derivative is applied to the raw waveform, considering the trade-off between efficiency and purity. Then, an integration is applied to the second derivative, and a threshold $T_{2}$ is required for peak detection. It is worth mentioning that only the positive part of the derivatives, which corresponding to the fast rising edge of the signals, are kept in the calculation. Because the large slope of the rising edge is the key feature to characterize the signal.
The derivative-based algorithm is a simple and easy-to-implement algorithm for cluster counting reconstruction. However, it only uses partial information of the waveform and is less efficient in high noise levels. For high noise-level environment, a low-pass digital filter can be applied before applying the derivative-based algorithm. However, this could drastically reduce the efficiency as the signals of cluster counting are highly piled-up.

### 11 Apr 2024 v2

@@ 1 Introduction
One major challenge of the cluster counting is the efficient reconstruction algorithm to recover the clusters signals (in the form of peaks) in the waveform from a DC measurement. These cluster signals tend to be stacked together in noisy environments (Fig. 1). Traditional methods often fall short of achieving the necessary efficiency due to sub-optimal information utilization. For instance, in the derivative-based method (section A.1), the peaks are detected by analyzing derivatives and imposing threshold criteria. However, only the rising edges of the single pulses, instead of the full pulse shape, are effectively used. The machine learning (ML) algorithms, which are designed to harness large datasets to reduce complexity and find new features in data, are the state-of-the-art in PID.
@@
Traditional ML relies on fully labeled samples for supervised learning. Such method is well-suited for Monte Carlo (MC) simulated samples, where the MC samples are generated through first-principle simulation, ensuring perfect labels. However, when applying ML algorithm to real experimental data samples, it is usually difficult to apply fully supervised learning. Usually more or less significant discrepancies exist between the simulation and the experimental data, which leads to a degraded performance of the supervised model trained solely on simulated sample. In some cases, these models may even fail to predict experimental data accurately. On the other hand, the experimental data often lack precise labels for fully supervised learning.
Domain adaptation (DA) aims at alleviating this issue by transferring knowledge between domains pan2009survey . In our context, the MC simulated sample serves as the source domain, while the experimental data represents target domain. A principled approach to solving domain adaptation involves aligning the source domain distribution with that of the target domain. This alignment allows us to leverage labeled data from the source domain to train a classifier in the aligned domain, which can then be applied to the target domain. In reference flamary2016optimal , the Optimal Transport (OT) is exploited to align the source and target domains. One important property of OT is the geometry sensitivity, which facilitates distance computation between probability distributions. By relaxing the invariant requirement on conditional distribution, reference courty2017joint proposed the Joint Distribution Optimal Transport (JDOT) , in which both marginal feature and conditional distributions are adapted by minimizing a global divergence between them. A further improvement on the path is to integrate the deep neural network and enable large scale learning damodaran2018deepjdot .
In the context of these studies, models are trained with a target sample without any labels, which is also known as the unsupervised domain adaptation. In our research, we build an approach upon the work presented in reference damodaran2018deepjdot and create a semi-supervised version of the domain adaptation model. The model will be employed for analyzing data collected during the drift-tube-prototype experiment conducted at CERN cern . We make our code available at https://github.com/littlepi/PeakFindingSemiDA.
@@ 2 Data Samples for Cluster Counting
The major mechanism underlying for particle identification in gaseous detector is the ionization of matter by charged particles. For DC, each electron produced through ionization generates a distinct peak in the readout waveform. If the electronics is sufficiently fast, these peaks become quantifiable. The reconstruction algorithm plays a crucial role in identifying and associating these peaks with individual electrons. The waveforms can be collected either from test beam experiment or from simulations. In the subsequent sections, we delve into the specifics of the collected data samples.
@@ 4.1 Experiment with Pseudo Data
Table 1: Area under curve (AUC) and partial-AUC results for models. The partial-AUC’s are the normalized AUC values with false positive rates (FPR) smaller than 0.1.
@@ 5 Summary
Cluster counting in drift chamber represents a significant advancement in PID techniques for particle physics experiments. The reconstruction algorithm is a critical challenge in this context. Our semi-supervised DA algorithm achieves state-of-the-art performance for peak finding in real experimental data. Importantly, the algorithm’s principles can be extended to other DA problems that exhibit geometric sensitivity. Looking ahead, as more test beam data become available across a wider range of beam momenta, our algorithm will play a crucial role in data analysis and may be implemented in FPGA systems for feasibility studies related to cluster counting.
@@ Acknowledgements
This research was funded by Institute of High Energy Physics (Chinese Academy of Sciences) Innovative Project on Sciences and Technologies under Contracts Nos. E3545BU210 and E25456U210, National Natural Science Foundation of China (NSFC) under Contract Nos. 12275296.
@@ A.1 Derivative-Based Algorithm
The principle of a derivative-based algorithm is intuitive. If a peak has a fast rising edge, the derivative should be larger than some threshold, and one can detect peaks by requiring threshold passing. The pseudo code of the algorithm is summarized in Algorithm 1. To exclude noise in the baseline, a threshold $T_{1}$ of approximately three times the RMS of the average noise amplitudes is initially required. Higher order derivatives are efficient for hidden-peak detection, but more sensitive to noise. Therefore, a second-order derivative is applied to the raw waveform, considering the trade-off between efficiency and purity. Then, an integration is applied to the second derivative, and a threshold $T_{2}$ is required for peak detection. Since large slope of the rising edge is a key feature to characterize the signal, only the positive derivatives corresponding to it are kept in the calculation.
The derivative-based algorithm is a simple and easy-to-implement algorithm for cluster counting reconstruction. However, it only uses partial information of the waveform and is less efficient in high noise levels. For high noise-level environment, a low-pass digital filter can be applied before applying the derivative-based algorithm. However, this could drastically reduce the efficiency as the signals of clusters are highly piled-up.